In SM Cabanatuan, Bench clothing department store advertises a 30% off sale. On that same day, an additional 10% will be given for the first 100 customers for the same brand.

Answer the following questions:

a. Assume that you are one of the first 100 lucky customers, use composition of function to find the total percent discount.

b. What would be the sale price of a PHP 900 jacket?
1
Expert's answer
2021-09-14T06:06:08-0400

a. Let "x=" the initial product price. Then

"f(x)=(1-0.3)x=0.7x,"

"g(x)=(1-0.1)x=0.9x"

"(f\circ g)(x)=0.7(0.9x)=0.63x"

The total percent discount is 37%.


b.


"0.63(900)=567"

The sale price of a PHP 900 jacket would be PHP 567.